1 Average minimum rank of a graph

By Francesco Barioli, Shaun Fallat, Tracy Hall, Daniel Hershkowitz, Leslie Hogben, Ryan Martin, Bryan Shader, Hein van der Holst.

We establish asymptotic upper and lower bounds on the average minimum rank of a graph using probabilistic, linear algebraic and graph theoretic techniques.